# The Effect of Personalization in FedProx: A Fine-grained Analysis on Statistical Accuracy and Communication Efficiency

### v3

@@ B.5 Proof of The Local Model Convergence Rate in Theorem 2
Appendix C Additional Experiment Details and Results
C.1 Additional Experiment Details
In this section, we provide more details on the implementation of the methods used in the empirical analysis.
@@ C.1 Additional Experiment Details
Real Dataset. The MNIST dataset consists of 70,000 grayscale images of handwritten digits, each 28x28 pixels, classified into 10 classes (digits 0-9). The EMNIST dataset extends MNIST, including handwritten letters and digits. The Balanced split contains 131,600 28x28 grayscale images across 47 balanced classes (digits and uppercase/lowercase letters). And the CIFAR-10 dataset contains 60,000 color images (32x32 pixels, 3 RGB channels), categorized into 10 classes representing real-world objects (e.g., airplane, car, and dog).
@@
Algorithms. To fully investigate the effect of personalization in FedProx with the Algorithm B.1, we compare the two extreme algorithms (GlobalTraining and LocalTraining). Furthermore, since there exist other methods solving the objective (1) to do personalized federated learning, we also try to investigate the effect of personalization in other methods, like pFedMe (T Dinh et al., 2020). The algorithm pFedMe has a similar design with a double loop to solve the subproblem in each communication round.
@@
Selection of $\lambda$. For the logistic regression on the synthetic dataset, we selected three values for $\lambda$: small $\lambda=0.02$, medium $\lambda=0.1$, and large $\lambda=0.5$. For logistic regression on the MNIST dataset, we used small $\lambda=0.5$, medium $\lambda=1.5$, and large $\lambda=2.5$. For CNN on the MNIST dataset, we used small $\lambda=0.2$, medium $\lambda=0.5$, and large $\lambda=2.5$. Such a choice is to make sure that each line in the figure is clearly separable. Additionally, we implemented an adaptive choice of $\lambda$ following the formula outlined in Theorem 1, with $\rho$ set to 2. The value of $\rho$ was determined through grid search over the range $[1,10]$ to ensure optimal performance. For other real datasets, the regularization term $\lambda$ is set as small $\lambda=0.1$, medium $\lambda=$ 0.5, and large $\lambda=$ 1.
@@
Selection of Step Size. For the synthetic dataset, aligned with Theorem 2, we set the global step size $\gamma=\frac{\lambda+L}{\lambda L}$ and the local step size $\eta=\frac{1}{L+\lambda}$. The smoothness constant $L$ is computed as the upper bound of the $L$-smoothness constant for logistic regression, specifically $4^{-1}\Lambda_{\text{max}}$, where $\Lambda_{\text{max}}$ denotes the largest eigenvalue of the Gram matrix $\bm{X}^{\top}\bm{X}$. The local step size for the synthetic dataset was further tuned through grid search in $\{0.001,0.01,0.1\}$. For the real dataset, the global learning rate was set to $\frac{1}{\lambda}$, while the local learning rate was set to 0.01, chosen via grid search. For both GlobalTrain and LocalTrain training baselines, the learning rate was also chosen via grid search within the range $\{0.001,0.01,0.1\}$. LocalTrain is implemented using a local (S)GD, and GlobalTrain is implemented using FedAvg (McMahan et al., 2017).
@@
Optimizer. For the synthetic dataset, following the analysis in the paper, the loss function is optimized using gradient descent (GD). For the real dataset, all models are implemented in PyTorch, and optimization is performed using the stochastic gradient descent (SGD) solver.

### v4

@@ B.3.2 Proof of The Local Model Convergence Rate in Theorem 2
B.4 Proof of Corollary 2
where we have used the fact that
$K\geq 2+\frac{\lambda+L}{\lambda+\mu}\cdot\log(1056\kappa^{2}).$ (134)
B.5 Global Convergence Rate with Stochastic Noise
Similar to the noiseless case, we propose FedCLUP below which uses stochastic gradient descent to solve Problem (2).
Algorithm 2 FedCLUP: Federated Learning with Constant Local Update Personalization
Input: Initial global model ${\bm{w}}_{1}^{(g)}$, initial local models $\{{\bm{w}}_{0,K}^{(i)}\}_{i\in[m]}$, global rounds $T$, global step sizes $\gamma$, local rounds $K$, local step sizes $\eta$, stochastic batch size $B$
for $t=1,\dots,T$ do
       The server sends ${\bm{w}}_{t}^{(g)}$ to client $i$, $\forall i\in[m]$;
Each client randomly draw $B$ data without replacement;
Set ${\bm{w}}_{t,0}^{(i)}={\bm{w}}_{t-1,K}^{(i)}$for $k=0,\dots,K-1$ do
             ${\bm{w}}_{t,k+1}^{(i)}={\bm{w}}_{t,k}^{(i)}-\frac{\eta}{B}\sum_{j\in[B]}\left\{\nabla\ell({\bm{w}}_{t,k}^{(i)},z_{i,j})+\lambda\left({\bm{w}}_{t,k}^{(i)}-\bm{w}_{t}^{(g)}\right)\right\}$
      Push ^$\widehat{\nabla}F_{i}(\bm{w}_{t}^{(g)})=\lambda(\bm{w}_{t}^{(g)}-{\bm{w}}_{t,K}^{(i)})$ to the server  $\bm{w}_{t+1}^{(g)}\leftarrow\bm{w}_{t}^{(g)}-\frac{\gamma}{m}\sum_{i\in[m]}\widehat{\nabla}F_{i}(\bm{w}_{t}^{(g)})$
We use the following assumption to characterize the stochastic noise when sampling new data in each iteration over all clients without replacement.
Assumption 4 (Stochastic noise).
For any $i\in[m],j\in[n_{i}],k\in[K],t>0$, we assume
$\mathbb{E}\left[\left\|\nabla_{\bm{w}_{i}}\hat{h}_{i}(\bm{w}_{t,k}^{(i)},\bm{w}_{t}^{(i)},s_{ij})-\nabla_{\bm{w}_{i}}h_{i}(\bm{w}_{t,\star}^{(i)},\bm{w}_{t}^{(i)})\right\|^{2}\right]\leq\sigma^{2}$ (135)
Let we define $\mathcal{F}_{t,k}$ to be the sigma algebra generated by the randomness by Algorithm 2 up to ${\bm{w}}_{t,k}^{(i)}$.
Lemma 7 (Inner Loop Error Contraction with stochastic).
$\begin{split}g_{t+1}=&\;\mathbb{E}\left(\left\|\frac{1}{m}\sum_{i\in[m]}\bm{w}_{t+1,k}^{(i)}-\bm{w}_{t+1,*}^{(i)}\right\|^{2}\Big{|}\mathcal{F}_{t+1,k}\right)\\ &\leq\left[1+\epsilon+2\left(1+\frac{1}{\epsilon}\right)L_{w}^{2}\gamma^{2}\lambda^{2}\right]\left(1-\eta_{t}\mu_{\ell}\right)^{k}g_{t}\\ &\quad+2\left(1+\frac{1}{\epsilon}\right)\left(1-\mu_{\ell}\eta_{t}\right)^{k}L_{w}^{2}\gamma^{2}\|\nabla F(\bm{w}_{t}^{(g)})\|^{2}\\ &\quad+\frac{2\eta_{t}}{\mu_{\ell}}\frac{\sigma^{2}}{Bm}.\end{split}$ (136)
Proof.
$\begin{split}g_{t+1}=&\;\mathbb{E}\left(\left\|\frac{1}{m}\sum_{i\in[m]}\bm{w}_{t+1,k}^{(i)}-\bm{w}_{t+1,*}^{(i)}\right\|^{2}\Big{|}\mathcal{F}_{t+1,k}\right)\\ =&\;\mathbb{E}\left(\left\|\frac{1}{m}\sum_{i\in[m]}\left[\bm{w}_{t+1,k-1}^{(i)}-\frac{\eta_{t}}{B}\sum_{j}\nabla_{w_{i}}\hat{h}_{i}(\bm{w}_{t+1,k-1}^{(i)},\bm{w}_{t+1}^{(i)},s_{ij})-\bm{w}_{t+1,*}^{(i)}\right]\right\|^{2}\Big{|}\mathcal{F}_{t+1,k}\right)\\ =&\;\mathbb{E}\left(\left\|\frac{1}{m}\sum_{i}\bm{w}_{t+1,k-1}^{(i)}-\bm{w}_{t+1,*}^{(i)}\right\|^{2}\Big{|}\mathcal{F}_{t+1,k}\right)\\ &+\eta_{t}^{2}\mathbb{E}\left(\left\|\frac{1}{mB}\sum_{i}\sum_{j}\nabla_{w_{i}}\hat{h}_{i}(\bm{w}_{t+1,k-1}^{(i)},\bm{w}_{t+1}^{(i)},s_{ij})\right\|^{2}\Big{|}\mathcal{F}_{t+1,k}\right)\\ &-2\eta_{t}\mathbb{E}\left[\left\langle\frac{1}{m}\sum_{i}\bm{w}_{t+1,k-1}^{(i)}-\bm{w}_{t+1,*}^{(i)},\frac{1}{Bm}\sum_{j}\nabla_{w_{i}}\hat{h}_{i}(\bm{w}_{t+1,k-1}^{(i)},\bm{w}_{t+1}^{(i)},s_{ij})\right\rangle\Big{|}\mathcal{F}_{t+1,k}\right].\end{split}$ (137)
To simplify the result, we shall note
$\mathbb{E}\left(\left\|\frac{1}{mB}\sum_{i}\sum_{j}\nabla_{\bm{w}_{i}}\hat{h}_{i}(\bm{w}_{t+1,k-1}^{(i)},\bm{w}_{t+1}^{(i)},s_{ij})\right\|^{2}\Big{|}\mathcal{F}_{t+1,k}\right)$
$=$ $\;\mathbb{E}\left(\left\|\frac{1}{mB}\sum_{i}\sum_{j}\nabla_{\bm{w}_{i}}\hat{h}_{i}(\bm{w}_{t+1,k-1}^{(i)},\bm{w}_{t+1}^{(i)},s_{ij})-\nabla_{\bm{w}_{i}}h_{i}(\bm{w}_{t+1,k-1}^{(i)},\bm{w}_{t+1}^{(i)})\right\|^{2}\Big{|}\mathcal{F}_{t+1,k}\right)$
$+\left\|\frac{1}{m}\sum_{i}\nabla_{\bm{w}_{i}}h_{i}(\bm{w}_{t+1,k-1}^{(i)},\bm{w}_{t+1}^{(i)})\right\|^{2}$
$\overset{(a)}{\leq}$ $\;\frac{\sigma^{2}}{mB}+\left\|\frac{1}{m}\sum_{i}\left(\nabla_{\bm{w}_{i}}h_{i}(\bm{w}_{t+1,k-1}^{(i)},\bm{w}_{t+1}^{(i)})-\nabla_{\bm{w}_{i}}h_{i}(\bm{w}_{t+1,*}^{(i)},\bm{w}_{t+1}^{(i)})\right)\right\|^{2}$
$\overset{(b)}{\leq}$ $\;\frac{\sigma^{2}}{mB}+2L_{\ell}\frac{1}{m}\sum_{i}\left(h_{i}(\bm{w}_{t+1,k-1}^{(i)},\bm{w}_{t+1}^{(i)})-h_{i}(\bm{w}_{t+1,*}^{(i)},\bm{w}_{t+1}^{(i)})\right).$ (138)
where step (a) stands due to Assumption 4 and the optimality of $\bm{w}_{t+1,\star}^{(i)}$, step (b) uses the Lemma 2.29 in (Garrigos & Gower, 2023). Moreover, for the third term in (137), it yields
$\begin{split}&\mathbb{E}\left[\left\langle\frac{1}{m}\sum_{i}\bm{w}_{t+1,k-1}^{(i)}-\bm{w}_{t+1,*}^{(i)},\frac{1}{Bm}\sum_{i,j}\nabla_{\bm{w}_{i}}\hat{h}_{i}(\bm{w}_{t+1,k-1}^{(i)},\bm{w}_{t+1}^{(i)},s_{j})\right\rangle\Big{|}\mathcal{F}_{t+1,k}\right]\\ =&\left\langle\frac{1}{m}\sum_{i}\bm{w}_{t+1,k-1}^{(i)}-\bm{w}_{t+1,*}^{(i)},\frac{1}{m}\sum_{i}\nabla_{\bm{w}_{i}}h_{i}(\bm{w}_{t+1,k-1}^{(i)},\bm{w}_{t+1}^{(i)})\right\rangle\\ \geq&\frac{1}{m}\sum_{i}\left(h_{i}(\bm{w}_{t+1,k-1}^{(i)},\bm{w}_{t+1}^{(i)})-h_{i}(\bm{w}_{t+1,*}^{(i)},\bm{w}_{t+1}^{(i)})\right)+\frac{\mu_{\ell}}{2}\left\|\frac{1}{m}\sum_{i}\bm{w}_{t+1,k-1}^{(i)}-\bm{w}_{t+1,*}^{(i)}\right\|^{2},\\ \end{split}$ (139)
where the last line comes from the strong convexity of $h_{i}$ in Assumption 2.
Plugging the results (B.5) and (139) into (137), we obtain
$\begin{split}&\mathbb{E}\left[\left\|\frac{1}{m}\sum_{i}\bm{w}_{t+1,k}^{(i)}-\bm{w}_{t+1,*}^{(i)}\right\|^{2}\Big{|}\mathcal{F}_{t+1,k}\right]\\ &\leq\left(1-\eta_{t}\mu_{\ell}\right)\left\|\frac{1}{m}\sum_{i}\bm{w}_{t+1,k-1}^{(i)}-\bm{w}_{t+1,*}^{(i)}\right\|^{2}\\ &+2\eta_{t}(L_{\ell}\eta_{t}-1)\left(\frac{1}{m}\sum_{i}\left(h_{i}(\bm{w}_{t+1,k-1}^{(i)},\bm{w}_{t+1}^{(i)})-h_{i}(\bm{w}_{t+1,*}^{(i)},\bm{w}_{t+1}^{(i)})\right)\right)\\ &+\eta_{t}^{2}\frac{\sigma^{2}}{Bm}.\end{split}$ (140)
If we assume $\eta_{t}\leq\frac{1}{L_{\ell}}$ we can simplify (140) as
$\begin{split}&\mathbb{E}\left[\left\|\frac{1}{m}\sum_{i}\bm{w}_{t+1,k}^{(i)}-\bm{w}_{t+1,*}^{(i)}\right\|^{2}\Big{|}\mathcal{F}_{t+1,k}\right]\\ &\leq\left(1-\mu_{\ell}\eta_{t}\right)\left\|\frac{1}{m}\sum_{i}\bm{w}_{t+1,k-1}^{(i)}-\bm{w}_{t+1,*}^{(i)}\right\|^{2}+\eta_{t}^{2}\frac{\sigma^{2}}{Bm}\\ &\leq\left(1-\mu_{\ell}\eta_{t}\right)^{k}\left\|\frac{1}{m}\sum_{i}\bm{w}_{t,k}^{(i)}-\bm{w}_{t,*}^{(i)}+\bm{w}_{t,*}^{(i)}-\bm{w}_{t+1,*}^{(i)}\right\|^{2}+\sum_{l=0}^{k-1}\left(1-\mu_{\ell}\eta_{t}\right)^{k-l}\eta_{t}^{2}\frac{\sigma^{2}}{Bm}\\ &\leq(1+\epsilon)\left(1-\mu_{\ell}\eta_{t}\right)^{k}g_{t}+\sum_{l=0}^{k-1}\left(1-\mu_{\ell}\eta_{t}\right)^{k-l}\eta_{t}^{2}\frac{\sigma^{2}}{Bm}+\left(1+\frac{1}{\epsilon}\right)\left(1-\mu_{\ell}\eta_{t}\right)^{k}L_{w}^{2}\gamma^{2}\left(2\|\nabla F(\bm{w}_{t}^{(g)})\|^{2}+2\lambda^{2}g_{t}\right).\end{split}$ (141)
Reorganizing all the terms in (141), we get
$\begin{split}&\mathbb{E}\left[\left\|\frac{1}{m}\sum_{i}\bm{w}_{t+1,k}^{(i)}-\bm{w}_{t+1,*}^{(i)}\right\|^{2}\Big{|}\mathcal{F}_{t+1,k}\right]\\ &\leq\left[1+\epsilon+2\left(1+\frac{1}{\epsilon}\right)L_{w}^{2}\gamma^{2}\lambda^{2}\right]\left(1-\eta_{t}\mu_{\ell}\right)^{k}g_{t}+2\left(1+\frac{1}{\epsilon}\right)\left(1-\mu_{\ell}\eta_{t}\right)^{k}L_{w}^{2}\gamma^{2}\|\nabla F(\bm{w}_{t}^{(g)})\|^{2}+\frac{2\eta_{t}}{\mu_{\ell}}\frac{\sigma^{2}}{Bm}.\end{split}$ (142)
Theorem 3.
Suppose Assumptions 1, 2 and 4 hold. Let $\{{\bm{w}}_{t}^{(g)}\}_{t\geq 0}$ and $\{{\bm{w}}_{t,K}^{(i)}\}_{t\geq 0}$ be the sequence generated by Algorithm 2 with $\gamma<1/L_{g}$, $\eta\leq 1/L_{\ell}$, and the inner loop iteration number satisfying
then ${\bm{w}}_{t}^{(g)}$ converges to $\widetilde{\bm{w}}^{(g)}$ linearly at rate $21-(\gamma\mu_{g})/2-(\gamma\mu_{g})^{2}/$ 2 within a neighbor of
$\frac{8}{\textit{min}\{\frac{1}{2}\gamma\mu_{g},\mu_{\ell}\eta\}}\frac{1}{\mu_{g}^{2}}\lambda^{2}\eta^{2}\frac{\sigma^{2}}{Bm}.$
Proof.
Analogy to the proof of Theorem 2, we define
$q:=\left(1+\epsilon+2\left(1+\frac{1}{\epsilon}\right)L_{w}^{2}\gamma^{2}\lambda^{2}\right)\left(1-\eta_{t}\mu_{\ell}\right)^{k}.$ (144)
If we assume
$q$ $\leq 1-\gamma\mu_{g}\leq 1-\gamma\mu_{g}+\gamma\epsilon_{g},\quad$ (145)
and define
$c_{1}=\frac{1}{1-\gamma\mu_{g}+\gamma\epsilon_{g}-q}\left(\frac{\gamma}{\epsilon_{g}}+\gamma^{2}(2+\zeta^{-1})\right)\lambda^{2},$ (146)
combining Lemma 7 and Lemma 6, it yields
$\begin{split}&\mathbb{E}\left[\left\|\bm{w}_{t+1}^{(g)}-\tilde{\bm{w}}^{(g)}\right\|^{2}\Big{|}\mathcal{F}_{t+1,k}\right]+c_{1}g_{t+1}\\ &\leq(1-\gamma\mu_{g}+\gamma\epsilon_{g})\left(\left\|\bm{w}^{(g)}_{t}-\tilde{\bm{w}}^{(g)}\right\|^{2}+c_{1}g_{t}\right)-\gamma^{2}(1-\gamma L_{g}(1+\zeta))\left\|\nabla F(\bm{w}_{t}^{(g)})\right\|^{2}\\ &\quad+\left(\frac{2}{1-\gamma\mu_{g}+\gamma\epsilon_{g}-q}\left(\frac{\gamma}{\epsilon_{g}}+\gamma^{2}(2+\zeta^{-1})\right)\lambda^{2}\right)2\left(1+\frac{1}{\epsilon}\right)\left(1-\mu_{\ell}\eta_{t}\right)^{k}L_{w}^{2}\gamma^{2}\left\|\nabla F(\bm{w}_{t}^{(g)})\right\|^{2}\\ &\quad+c_{1}\frac{2\eta_{t}}{\mu_{\ell}}\frac{\sigma^{2}}{Bm}.\end{split}$
For $\left\|\nabla F(\bm{w}_{t}^{(g)})\right\|^{2}$, it suffices to show
$\frac{4\gamma}{1-\gamma\mu_{g}+\gamma\epsilon_{g}-q}\left(\frac{1}{\epsilon_{g}}+\gamma(2+\zeta^{-1})\right)\lambda^{2}(1+\frac{1}{\epsilon})\left(1-\mu_{\ell}\eta_{t}\right)^{k}L_{w}^{2}\leq 1-\gamma L_{g}(1+\zeta).$
If we set $\epsilon_{g}=\frac{\mu_{g}}{2}$, it yields the convergence rate
$1-\frac{\gamma\mu_{g}}{2}.$ (148)
And the condition (B.5) can be reformulated as
$\frac{4}{\epsilon_{g}}\left(\frac{1}{\epsilon_{g}}+\gamma(2+\zeta^{-1})\right)\lambda^{2}(1+\epsilon^{-1})\left(1-\mu_{\ell}\eta_{t}\right)^{k}L_{w}^{2}\leq 1-\gamma L_{g}(1+\zeta).$
Restrict $\gamma$
$\gamma:$ $\quad\gamma(2+\zeta^{-1})\leq\frac{1}{\epsilon_{g}}.$ (150)
Then condition (B.5) yields
$\frac{8}{\epsilon_{g}^{2}}\lambda^{2}(1+\frac{1}{\epsilon})\left(1-\mu_{\ell}\eta_{t}\right)^{k}L_{w}^{2}\leq 1-\gamma L_{g}(1+\zeta).$ (151)
Then if we assume $\epsilon=1$ and $\zeta=1-\gamma L_{g}$, the conditions (145), (150) and (151) can be satisfied with the following sufficient conditions
$\begin{split}&\gamma<\frac{1}{L_{g}},\\ &\eta_{t}\leq\frac{1}{L_{\ell}},\\ &(2+4L_{w}^{2}\gamma^{2}\lambda^{2})\left(1-\eta_{t}\mu_{\ell}\right)^{k}\leq 1-\gamma\mu_{g},\\ &\frac{64}{\mu_{g}^{2}}\left(1-\eta_{t}\mu_{\ell}\right)^{k}L_{w}^{2}\lambda^{2}\leq(1-\gamma L_{g})^{2}.\end{split}$ (152)
And we can get a more straightforward sufficient condition
$\gamma<\frac{1}{L_{g}},$ (153)
$\eta_{t}\leq\frac{1}{L_{\ell}},$ (154)
$\left(2+64\frac{L_{w}^{2}}{\mu_{g}^{2}}\lambda^{2}\right)\left(1-\eta_{t}\mu_{\ell}\right)^{k}\leq(1-\gamma L_{g})^{2}.$ (155)
Under such a sufficient condition, the contraction of the combination of inner loop and the outer loop would be rewritten as
$\begin{split}&\mathbb{E}\left[\left\|\bm{w}_{t+1}^{(g)}-\tilde{\bm{w}}^{(g)}\right\|^{2}\Big{|}\mathcal{F}_{t+1,k}\right]+c_{1}g_{t+1}\\ &\leq(1-\gamma\mu_{g}+\gamma\epsilon_{g})\left(\left\|\bm{w}^{(g)}_{t}-\tilde{\bm{w}}^{(g)}\right\|^{2}+c_{1}g_{t}\right)\quad+\sum_{l=0}^{t}\left(1-\frac{1}{2}\gamma\mu_{g}\right)^{t-l}\frac{2}{\epsilon_{g}^{2}}\lambda^{2}\sum_{l=0}^{k-1}\left(1-\mu_{\ell}\eta\right)^{k-l}\eta^{2}\frac{\sigma^{2}}{Bm}\\ &\leq(1-\frac{1}{2}\gamma\mu_{g})^{t+1}\left(\left\|\bm{w}^{(g)}_{0}-\tilde{\bm{w}}^{(g)}\right\|^{2}+c_{1}g_{0}\right)+\sum_{l=0}^{t}\left(1-\frac{1}{2}\gamma\mu_{g}\right)^{t-l}\frac{2}{\epsilon_{g}^{2}}\lambda^{2}\sum_{l=0}^{k-1}\left(1-\mu_{\ell}\eta\right)^{k-l}\eta^{2}\frac{\sigma^{2}}{Bm}\\ &\leq(1-\frac{1}{2}\gamma\mu_{g})^{t+1}\left(\left\|\bm{w}^{(g)}_{0}-\tilde{\bm{w}}^{(g)}\right\|^{2}+c_{1}g_{0}\right)+\sum_{l=0}^{tk-1}2\left(\text{max}(1-\frac{1}{2}\gamma\mu_{g},1-\mu_{\ell}\eta)\right)^{l}\frac{1}{\epsilon_{g}^{2}}\lambda^{2}\eta^{2}\frac{\sigma^{2}}{Bm}\\ &\leq(1-\frac{1}{2}\gamma\mu_{g})^{t+1}\left(\left\|\bm{w}^{(g)}_{0}-\tilde{\bm{w}}^{(g)}\right\|^{2}+c_{1}g_{0}\right)+\frac{8}{\text{min}\{\frac{1}{2}\gamma\mu_{g},\mu_{\ell}\eta\}}\frac{1}{\mu_{g}^{2}}\lambda^{2}\eta^{2}\frac{\sigma^{2}}{Bm}.\end{split}$ (156)
With $\eta=\text{min}\{\frac{\mu(\lambda+L)}{4L(\lambda+\mu)^{2}},\frac{1}{\lambda+L}\}$, $\gamma=1/(2L_{g})=\frac{\lambda+L}{2\lambda L}$, we can get the solution within an error ball $\left(\mathcal{O}(\varepsilon)+\text{min}\{\frac{1}{4\mu L},\frac{\lambda+\mu}{\mu^{2}(\lambda+L)}\}\frac{\sigma^{2}}{Bm}\right)$ with the communication cost $\mathcal{O}\big{(}\kappa_{g}\log(1/\varepsilon)\big{)}=\mathcal{O}\left(\kappa\cdot\frac{\lambda+\mu}{\lambda+L}\cdot\log\frac{1}{\varepsilon}\right)$ and the computation cost $\tilde{\mathcal{O}}\left(\kappa\cdot\log\frac{1}{\varepsilon}\right)$.
As stated in Theorem 4.2 of (Hanzely & Richtárik, 2020), when $p^{\star}=\frac{\lambda}{\lambda+L}$, the stochastic noise term is independent of the number of clients. Consequently, it does not theoretically demonstrate the expected benefit of collaboration in reducing noise. In contrast, our analysis reveals that even when adjusting $\lambda$, the noise remains of the same order, specifically $\mathcal{O}\left(\frac{1}{Bm}\right)$.
Appendix C Proof in Section 5
Corollary 4.
Under the assumptions of Theorem 1 and Corollary 2, ${\bm{w}}_{T,K}^{(i)}$ generated by Algorithm 1 satisfies
$\begin{split}&\mathbb{E}\left\|{\bm{w}}_{T,K}^{(i)}-{{\bm{w}}_{\star}^{(i)}}\right\|^{2}\\ &\leq\left(1-\frac{1}{\kappa}-\frac{L-\mu}{\kappa(\lambda+\mu)}\right)^{T}\mathbb{E}\left\|{\bm{w}}_{0,0}^{(i)}-\widetilde{\bm{w}}^{(i)}\right\|^{2}\\ &+\mathcal{O}\left[\left(\frac{1}{N}+R^{2}+\frac{1}{q_{1}(\lambda)}\right)\wedge\left(\frac{1}{n}+q_{2}(\lambda)\right)\right]\end{split}$ (157)
for any $i\in[m]$, where $q_{1}(\lambda)$ and $q_{2}(\lambda)$ are monotone increasing functions of $\lambda$ with $\lim_{\lambda\to+\infty}q_{1}(\lambda)=+\infty$ and $\lim_{\lambda\to 0}q_{2}(\lambda)=0$.
Proof.
The corollary is a direct result of Theorem 1 and Theorem 2, where $q_{1}(\lambda)$ and $q_{2}(\lambda)$ are defined in (70) and (67).
Appendix D Additional Experiment Details and Results
D.1 More Experiment Details
Federated Learning (FL) has emerged as a widely adopted approach with numerous real-world applications, like Large Language Model (LLM) (Bai et al., 2024; Wu et al., 2024; Zhao et al., 2023), Reinforcement Learning (Zheng et al., 2024, 2025; Zhang et al., 2025), Diffusion Model (Li et al., 2024a). In this section, we provide more details on the implementation of the methods used in the empirical analysis.
@@ D.1 More Experiment Details
Real Dataset. The MNIST dataset consists of 70,000 grayscale images of handwritten digits, each 28x28 pixels, classified into 10 classes (digits 0-9). The EMNIST dataset extends MNIST, including handwritten letters and digits. The Balanced split contains 131,600 28x28 grayscale images across 47 balanced classes (digits and uppercase/lowercase letters). The CIFAR-10 dataset contains 60,000 color images (32x32 pixels, 3 RGB channels), categorized into 10 classes representing real-world objects (e.g., airplane, car, and dog).
@@
Algorithms. To fully investigate the effect of personalization in Problem (2) with the Algorithm B.1, we compare the two extreme algorithms (GlobalTraining and LocalTraining). Furthermore, since there exist other methods solving Problem (2) to do personalized federated learning, we also try to investigate the effect of personalization in other methods, like pFedMe (T Dinh et al., 2020). The algorithm pFedMe has a similar design with a double loop to solve the subproblem in each communication round.
@@
Selection of $\lambda$. For the logistic regression on the synthetic dataset, we selected three values for $\lambda$: small $\lambda=0.02$, medium $\lambda=0.1$, and large $\lambda=0.5$. For logistic regression on the MNIST dataset, we used small $\lambda=0.5$, medium $\lambda=1.5$, and large $\lambda=2.5$. For CNN on the MNIST dataset, we used small $\lambda=0.2$, medium $\lambda=0.5$, and large $\lambda=2.5$. Such a choice is to make sure that each line in the figure is clearly separable. Additionally, we implemented an adaptive choice of $\lambda$ following the formula outlined in Corrollary 1, with $\rho$ set to 2. The value of $\rho$ was determined through grid search over the range $[1,10]$ to ensure optimal performance. For other real datasets, the regularization term $\lambda$ is set as small $\lambda=0.1$, medium $\lambda=$ 0.5, and large $\lambda=$ 1.
@@
Selection of Step Size. For the synthetic dataset, aligned with Theorem 2, we set the global step size $\gamma=(\lambda+L)/(\lambda L)$ and the local step size $\eta=1/(L+\lambda)$. The smoothness constant $L$ is computed as the upper bound of the $L$-smoothness constant for logistic regression, specifically $4^{-1}\Lambda_{\text{max}}$, where $\Lambda_{\text{max}}$ denotes the largest eigenvalue of the Gram matrix $\bm{X}^{\top}\bm{X}$. The local step size for the synthetic dataset was further tuned through grid search in $\{0.001,0.01,0.1\}$. For the real dataset, the global learning rate was set to $1/\lambda$, while the local learning rate was set to 0.01, chosen via grid search. For both GlobalTrain and LocalTrain training baselines, the learning rate was also chosen via grid search within the range $\{0.001,0.01,0.1\}$. LocalTrain is implemented using a local (S)GD, and GlobalTrain is implemented using FedAvg (McMahan et al., 2017).
@@
Optimizer. For the synthetic dataset, following the analysis in the paper, the loss function is optimized using gradient descent (GD). For the MNIST dataset, all models are implemented in PyTorch, and optimization is performed using the stochastic gradient descent (SGD) solver.